Consider a bond with a 7% annual coupon and a face value of $1,200 Complete the following table. (Enter your responses rounded to two decimal places) Years to Yield to Current Maturity Maturity Price 5% 7% 7% 5%
Answer #1
Years to maturity yield to maturity Current price
3 .05 1266.976

3

.07 816.7300
4 .07 1200
6 .05 1325.217
6 .09 1086.61

coupon Rate、%(finn!al), face Ulues $1200 30 Coupon ment Coupon far Now. Yeas to Matuaty FMarker aice Yield to Malunty -o,5% , Yeos to Matuaity 3 3600 +3 C 180 t .150 = 2904-21 ) てるsat1200-8x 2 3600 18 c ·04: | 84 t 1200-C 1900 +○ 一 ( 1536- 14-1536-C 1-14C=1368 , p 005) ·05-184 t 1200-41/12oote)-(1404-4) xf.4 +1200-grec)う 褜) y,) .oa- 1200+C う 5+c 0 1200 + C う. 1.24c = 1380